OpenStudy (anonymous):

In a circuit that has a voltage of 120 V and a current of 5 A, what wattage will the circuit hold?

OpenStudy (anonymous):

is wattage is power?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if it is power then power = V * I that means 120*5 = 600watts

OpenStudy (anonymous):

thank you:)

OpenStudy (anonymous):

power =I*V